# The dual complex of a $G$-variety

### v1

@@ 3. Applications
Such a divisor symbol is incompressible roughly if it cannot arise as the symbol for an exceptional divisor of a blowup of a lower dimensional subvariety (for a precise definition, see [12, Definition 3.3]). Assuming by induction the nonlinearizability of the residual action on $k(Y)$, one can prove the incompressibility of the corresponding divisor symbol. It follows that such a symbol cannot appear in the invariant associated to a linearizable action of $G$ on $\mathbb{P}^{n}$.
For every pair $n,d$ in the theorem, there are smooth hypersurfaces to which the result applies. For instance, if
$X\coloneqq\left\{\sum_{i=0}^{n+1}a_{i}x_{i}^{d}=0\right\}\subset\mathbb{P}^{n+1},$
then $X$ is invariant under any subgroup of the action by multiplication by $d$th roots of unity on the coordinates. That group has rank $n+1$.
@@
Besides the groups in Theorem 3.2, some other interesting cases of 3.5 in low dimensions are known. For instance, if $G\cong{\mathbb{Z}}/3$ is the group acting by third roots of unity on the first coordinate $x_{0}$ in $\mathbb{P}^{3}$, then the induced action on any $G$-invariant smooth cubic surface $X$ is not linearizable, even though $X$ is rational. Indeed, a cyclic group of prime order in the Cremona group $\mathrm{Cr}(2)$ fixing a curve of positive genus is not even stably $G$-birational to a linear action on $\mathbb{P}^{2}$ [4, Corollary 1.2]. However, 3.5 appears to be open for low rank abelian groups acting on hypersurfaces of small degree and large dimension.

### v2

@@ 3. Applications
Such a divisor symbol is incompressible roughly if it cannot arise as the symbol for an exceptional divisor of a blowup of a lower dimensional subvariety (for a precise definition, see [15, Definition 3.3]). Assuming by induction the nonlinearizability of the residual action on $k(Y)$, one can prove the incompressibility of the corresponding divisor symbol. It follows that such a symbol cannot appear in the invariant associated to a linearizable action of $G$ on $\mathbb{P}^{n}$.
@@
Besides the groups in Theorem 3.2, some other interesting cases of 3.7 in low dimensions are known. For instance, if $G\cong{\mathbb{Z}}/3$ is the group acting by third roots of unity on the first coordinate $x_{0}$ in $\mathbb{P}^{3}$, then the induced action on any $G$-invariant smooth cubic surface $X$ is not linearizable, even though $X$ is rational. Indeed, a cyclic group of prime order in the Cremona group $\mathrm{Cr}(2)$ fixing a curve of positive genus is not even stably $G$-birational to a linear action on $\mathbb{P}^{2}$ [5, Corollary 1.2]. However, 3.7 appears to be open for low rank abelian groups acting on hypersurfaces of small degree and large dimension.